# Generative Adversarial Networks for Image Super-Resolution: A Survey

### v4

@@ Abstract.
Single image super-resolution (SISR) has played an important role in the field of image processing. Recent generative adversarial networks (GANs) can achieve excellent results on low-resolution images with small samples. However, there are little literatures summarizing different GANs in SISR. In this paper, we conduct a comparative study of GANs from different perspectives. We first take a look at developments of GANs. Second, we present popular architectures for GANs in big and small samples for image applications. Then, we analyze motivations, implementations and differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners, where these GANs are analyzed via integrating different network architectures, prior knowledge, loss functions and multiple tasks. Next, we compare performance of these popular GANs on public datasets via quantitative and qualitative analysis in SISR. Finally, we highlight challenges of GANs and potential research points for SISR.
@@
ccs: General and reference Surveys and overviews
@@ 1. Introduction
Single image super-resolution (SISR) is an important branch in the field of image processing (Wang et al., 2020a). It also aims to recover a high-resolution (HR) image over a low-resolution (LR) image (Yang et al., 2019), leading to in its wide applications in medical diagnosis (Isaac and Kulkarni, 2015), video surveillance (Zhang et al., 2010) and disaster relief (Zha et al., 2015) etc. For instance, in the medical field, obtaining higher-quality images can help doctors accurately detect diseases (Huang et al., 2017). Thus, studying SISR is very meaningful to academia and industry.
@@
To address SISR problem, researchers have developed a variety of methods based on degradation models of low-level vision tasks (Zhou et al., 2014). There are three categories for SISR in general, i.e., image itself information, prior knowledge and machine learning. In the image itself information, directly amplifying resolutions of all pixels in a LR image through an interpolation way to obtain a HR image was a simple and efficient method in SISR (Park et al., 2003), i.e., nearest neighbor interpolation (Rukundo and Cao, 2012), bilinear interpolation (Li and Orchard, 2001) and bicubic interpolation (Keys, 1981), etc. It is noted that in these interpolation methods, high-frequency information is lost in the up-sampling process (Park et al., 2003). Alternatively, reconstruction-based methods were developed for SISR, according to optimization methods (Hardeep et al., 2013). That is, mapping a projection into a convex set to estimate the registration parameters can restore more details of SISR (Stark and Oskoui, 1989). Although the mentioned methods can overcome the drawbacks of image itself information methods, they still suffered the following challenges: non-unique solution, slow convergence speed and higher computational costs. To prevent this phenomenon, the priori knowledge and image itself information were integrated into a frame to find an optimal solution to improve the quality of the predicted SR images (Elad and Feuer, 1997; Irani and Peleg, 1991). Using maximum a posteriori (MAP) can regularize a loss function to obtain a maximum probability for improving the efficiency (Cheeseman et al., 1996). Besides, machine learning methods can be presented to deal with SISR, according to relation of data distribution (Sun et al., 2008). On the basis of ensuring the image SR effect, sparse-neighbor-embedding-based (SpNE) method via partition the training data set into a set of subsets to accelerate the speed of SR reconstruction (Gao et al., 2012). There are also many other SR methods (Sun et al., 2008; Yan et al., 2015) that often adopt sophisticated prior knowledge to restrict the possible solution space with an advantage of generating flexible and sharp detail. However, the performance of these methods rapidly degrades when the scale factor is increased, and these methods tend to be time-consuming (Nasrollahi and Moeslund, 2014).
@@
To obtain a better and more efficient SR model, a variety of deep learning methods were applied to a large-scale image dataset to solve the super-resolution tasks. For instance, Dong et al. proposed a super-resolution convolutional neural network (SRCNN) based pixel mapping that used only three layers to obtain stronger learning ability than these of some popular machine learning methods on image super-resolution (Dong et al., 2015). Although the SRCNN had a good SR effect, it still faced problems in terms of shallow architecture and high complexity. To overcome challenges of shallow architectures, Kim et al. (Kim et al., 2016) designed a deep architecture by stacking some small convolutions to improve performance of image super-resolution. Tai et al. (Tai et al., 2017) relied on recursive and residual operations in a deep network to enhance learning ability of a SR model. To further improve the SR effect, Lee et al. (Lim et al., 2017) used weights to adjust residual blocks to achieve better SR performance. To extract robust information, the combination of traditional machine learning methods and deep networks can restore more detailed information for SISR (Wang et al., 2015). For instance, Wang et al. (Wang et al., 2015) embedded sparse coding method into a deep neural network to make a tradeoff between performance and efficiency in SISR. To reduce the complexity, an up-sampling operation is used in a deep layer in a deep CNN to increase the resolution of low-frequency features and produce high-quality images (Dong et al., 2016). For example, Dong et al. (Dong et al., 2016) directly exploited the given low-resolution images to train a SR model for improving training efficiency, where the SR network used a deconvolution layer to reconstructing HR images. There are also other effective SR methods. For example, Lai et al. (Lai et al., 2017) used Laplacian pyramid technique into a deep network in shared parameters to accelerate the training speed for SISR. Zhang et al. (Zhang et al., 2018a) guided a CNN by attention mechanisms to extract salient features for improving the performance and visual effects in image SISR.
Although the mentioned SR methods have obtained excellent effect in SISR, the obtained damaged images are insufficient in the real world, which limits the application of the above SR methods on real cameras. To address problem of small samples, generative adversarial nets (GANs) used generator and discriminator in a game-like manner to obtain good performance on image applications (Goodfellow et al., 2014; Wang et al., 2021c). Specifically, the generator can generate new samples, according existing samples (Gui et al., 2021). The discriminator is used to distinguish the samples from generator (Gui et al., 2021). Due to their strong learning abilities, GANs become popular image super-resolution methods (Bell-Kligler et al., 2019). However, there are few studies summarizing these GANs for SISR. Also, differing from previous work based deep learning techniques for image super-resolution, i.e., Refs.(Jiang et al., 2021; Anwar et al., 2020), we can not only refer to importance of GANs for low- and high-levels in terms of big and small samples, but also first deeply product a summary of GANs in image super-resolution, according to combination of different training ways (i.e., supervised, semi-supervised and unsupervised manners), network architectures, prior knowledge, loss functions and multiple tasks, which can makes readers easier know principle, improvements, superiority and inferiority of different GANs for image super-resolution. That is, in this paper, we conduct a comprehensive overview of 197 papers to show their performance, pros, cons, complexity, challenges and potential research points, etc. First, we show the effects of GANs for image applications. Second, we present popular architectures for GANs in big and small samples for image applications. Third, we analyze motivations, implementations and differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners, where these GANs are worked by combining different network architectures, prior knowledge, loss functions and multiple tasks for image super-resolution. Fifth, we compare these GANs using experimental setting, quantitative analysis (i.e., PSNR, SSIM, complexity and running time) and qualitative analysis. Finally, we report on potential research points and existing challenges of GANs for image super-resolution. The overall architecture of this paper is shown in Fig. 1.
The remainder of this survey is organized as follows. Section 2 resents the developments of GANs. Section 3 gives a brief introduction of basic GANs for image processing tasks. Section 4 focuses on introduction of existing GANs via three ways on SISR. Section 5 compares performance of mentioned GANs from Section 4 for SISR. Section 6 offers potential directions and challenges of GANs in image super-resolution. Section 7 concludes the overview.
@@
[…]
The main structure of this paper.
@@ 2. Developments of GANs
Traditional machine learning methods prefer to use prior knowledge to improve performance of image processing applications (Sun et al., 2010). For instance, Sun et al. (Sun et al., 2010) proposed a gradient profile to restore more detailed information for improving performance of image super-resolution. Although machine learning methods based prior knowledge has fast execution speed, they have some drawbacks. First, they required manual setting parameters to achieve better performance on image tasks. Second, they required complex optimization methods to find optimized parameters. According to mentioned challenges, deep learning methods are developed (Bird et al., 2022). Deep learning methods used deep networks, i.e., CNNs to automatically learn features rather than manual setting parameters to obtain effective effects in image processing tasks, i.e., image classification (Bird et al., 2022), image inpainting (Liu et al., 2021) and image super-resolution (Wang et al., 2020a). Although these methods are effective big samples, they are limited for image tasks with small samples (Goodfellow et al., 2014).
To address problems above, GANs are presented in image processing (Goodfellow et al., 2014; Saxena and Cao, 2021). GANs consist of generator network and discriminator network. The generator network is used to generate new samples, according to given samples. The discriminator network is used to determine truth of obtained new samples. When generator and discriminator is balance, a GAN model is finished. The work process of GAN can be shown in Fig. 2, where G and D denote a generator network and discriminator network. To better understand GANs, we introduce several basic GANs as follows.
@@
[…]
The original structure of GAN.
To obtain more realistic effects, conditional information is fused into a GAN (CGAN) to randomly generate images, which are closer to real images (Mirza and Osindero, 2014). CGAN improves GAN to obtain more robust data, which has an important reference value to GANs for computer vision applications. Subsequently, increasing the depth of GAN instead of the original multilayer perceptron in a CNN to improve expressive ability of GAN is developed for complex vision tasks (Radford et al., 2015). To mine more useful information, the bidirectional generative adversarial network (BiGAN) used dual encoders to collaborate a generator and discriminator to obtain richer information for improving performance in anomaly detection, which is shown in Fig. 3 (Donahue et al., 2016). In Fig. 3, x denotes a feature vector, E is an encoder and y expresses an image from discriminator.
@@
[…]
The structure of BiGAN.
It is known that pretrained operations can be used to accelerate the training speed of CNNs for image recognition (Hinterstoisser et al., 2018). This idea can be treated as an energy drive. Inspired by that, Zhao et al. proposed an energy-based generative adversarial network (EBGAN) by using a pretraining operation into a discriminator to improve the performance in image recognition (Zhao et al., 2016). To keep consistency of obtained features with original images, cycle-consistent adversarial network (CycleGAN) relies on a cyclic architecture to achieve an excellent style transfer effect (Zhu et al., 2017) as illustrated in Fig. 4.
@@
[…]
The structure of CycleGAN.
Although pretrained operations are useful for training efficiency of network models, they may suffer from mode collapse. To address this problem, Wasserstein GAN (WGAN) used weight clipping to enhance importance of Lipschitz constraint to improve the stability of training a GAN (Arjovsky et al., 2017). WGAN used weight clipping to perform well. However, it is easier to cause gradient vanishing or gradient exploding (Gulrajani et al., 2017). To resolve this issue, WGAN used a gradient penalty (treated as WGAN-GP) to break the limitation of Lipschitz for pursuing good performance in computer vision applications (Brock et al., 2018). To further improve results of image generation, GAN enlarged batch size and used truncation trick as well as BIGGAN can make a tradeoff between variety and fidelity (Brock et al., 2018). To better obtained features of different parts of an image (i.e., freckles and hair), style-based GAN (StyleGAN) uses feature decoupling to control different features and finish style transfer for image generation (Karras et al., 2019). The architecture of StyleGAN and its generator are shown in Fig. 5 and Fig. 6.
@@ 3. Popular GANs for image applications
According to mentioned illustrations, it is known that variants of GANs based on properties of vision tasks are developed in Section 2. To further know GANs, we show different GANs on training data, i.e., big samples and small samples for different high- and low-level computer vision tasks as shown in Fig. 7.
Figure 7. Frame of popular GANs for image applications.
[…]
The structure of the Section 3.
@@ 3.1.1. GANs on big samples for image generation
Good performance of image generation depends on rich samples. Inspired by that, GANs are improved for image generation (Gui et al., 2021). That is, GANs use generator to produce more samples from high-dimensional data to cooperate discriminator for promoting results of image generation. For instance, boundary equilibrium generative adversarial networks (BEGAN) used obtained loss from Wasserstein to match loss of auto-encoder in the discriminator and achieve a balance between a generator and discriminator, which can obtain more texture information than that of common GANs in image generation (Berthelot et al., 2017). To control different parts of a face, StyleGAN decoupled different features to form a feature space for finishing transfer of texture information (Karras et al., 2019). Besides, texture synthesis is another important application of image generation (Li and Wand, 2016). For instance, Markovian GANs (MGAN) can quickly capture texture date of Markovian patches to achieve function of real-time texture synthesis (Li and Wand, 2016), where Markovian patches can be obtained Ref. (Gui et al., 2021). Periodic spatial GAN (PSGAN) (Bergmann et al., 2017) is a variant of spatial GAN (SGAN) (Jetchev et al., 2016), which can learn periodic textures of big datasets and a single image. These methods can be summarized in Table 2.

### v5

@@ Abstract.
Single image super-resolution (SISR) has played an important role in the field of image processing. Recent generative adversarial networks (GANs) can achieve excellent results on low-resolution images. However, there are little literatures summarizing different GANs in SISR. In this paper, we conduct a comparative study of GANs from different perspectives. We begin by surveying the development of GANs and popular GAN variants for image-related applications, and then analyze motivations, implementations and differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners, where these GANs are analyzed via integrating different network architectures, prior knowledge, loss functions and multiple tasks. Secondly, we compare the performances of these popular GANs on public datasets via quantitative and qualitative analysis in SISR. Finally, we highlight challenges of GANs and potential research points for SISR.
@@
$c$s: General and reference Surveys and overviews
@@ 1. Introduction
Single image super-resolution (SISR) is an important branch in the field of image processing (Wang et al., 2020b). It aims to recover a high-resolution (HR) image over a low-resolution (LR) image (Yang et al., 2019), leading to its wide applications in medical diagnosis (Isaac and Kulkarni, 2015), video surveillance (Zhang et al., 2010) and disaster relief (Zha et al., 2015) etc. For instance, in the medical field, obtaining higher-quality images can help doctors accurately detect diseases (Huang et al., 2017). Thus, studying SISR is very meaningful to academia and industry.
@@
To address SISR problem, researchers have developed a variety of methods based on degradation models of low-level vision tasks (Zhou et al., 2014). There are three categories for SISR in general, i.e., image itself information, prior knowledge and machine learning. In the image itself information, directly amplifying resolutions of all pixels in a LR image through an interpolation way to obtain a HR image was a simple and efficient method in SISR (Park et al., 2003), i.e., nearest neighbor interpolation (Rukundo and Cao, 2012), bilinear interpolation (Li and Orchard, 2001) and bicubic interpolation (Keys, 1981), etc. It is noted that in these interpolation methods, high-frequency information is lost in the up-sampling process (Park et al., 2003), which may decrease performance in image super-resolution. Alternatively, reconstruction-based methods were developed for SISR, according to optimization methods (Hardeep et al., 2013). That is, mapping a projection into a convex set to estimate the registration parameters can restore more details of SISR (Stark and Oskoui, 1989). Although the mentioned methods can overcome the drawbacks of image itself information methods, they still suffered from the following challenges: non-unique solution, slow convergence speed and higher computational costs. To prevent this phenomenon, the prior knowledge and image itself information were integrated into a frame to find an optimal solution to improve the quality of the predicted SR images (Elad and Feuer, 1997; Irani and Peleg, 1991). Besides, machine learning methods can be presented to deal with SISR, according to relation of data distribution (Sun et al., 2008). There are also many other SR methods (Sun et al., 2008; Yan et al., 2015) that often adopt sophisticated prior knowledge to restrict the possible solution space with the advantage of generating flexible and sharp detail. However, the performance of these methods rapidly degrades when the scale factor is increased, and these methods tend to be time-consuming (Nasrollahi and Moeslund, 2014).
@@
To obtain a better and more efficient SR model, a variety of deep learning methods were applied to a large-scale image dataset to solve the super-resolution tasks (Guo et al., 2023; Gao et al., 2023a). For instance, Dong et al. proposed a super-resolution convolutional neural network (SRCNN) based pixel mapping that used only three layers to obtain stronger learning ability than these of some popular machine learning methods on image super-resolution (Dong et al., 2015). Although the SRCNN had a good SR effect, it still faced problems in terms of shallow architecture and high complexity. To overcome challenges of shallow architectures, Kim et al. (Kim et al., 2016) designed a deep architecture by stacking some small convolutions to improve performance of image super-resolution. Tai et al. (Tai et al., 2017) relied on recursive and residual operations in a deep network to enhance learning ability of a SR model. To further improve the SR effect, Lee et al. (Lim et al., 2017) used weights to adjust residual blocks to achieve better SR performance. To extract robust information, the combination of traditional machine learning methods and deep networks can restore more detailed information for SISR (Wang et al., 2015). For instance, Wang et al. (Wang et al., 2015) embedded sparse coding method into a deep neural network to make a tradeoff between performance and efficiency in SISR. To reduce the complexity, an up-sampling operation is used in a deep layer in a deep CNN to increase the resolution of low-frequency features and produce high-quality images (Dong et al., 2016). For example, Dong et al. (Dong et al., 2016) directly exploited the given low-resolution images to train a SR model for improving training efficiency, where the SR network used a deconvolution layer to reconstruct HR images. There are also other effective SR methods. For example, Lai et al. (Lai et al., 2017) used Laplacian pyramid technique into a deep network in shared parameters to accelerate the training speed for SISR. Zhang et al. (Zhang et al., 2018b) guided a CNN by attention mechanisms to extract salient features for improving the performance and visual effects in image SISR. Dynamic super resolution network (DSRNet) (Tian et al., 2024b) proposes a dynamic network architecture that utilizes a dynamic gate mechanism to dynamically adjust network parameters to adapt to different scenarios, significantly enhancing the robustness and applicability of the image super-resolution model in complex scenes. Tian et al. (Tian et al., ) proposed a cosine network-based method for SISR, which introduced odd and even enhancement blocks to extract complementary homologous structural information and incorporated a cosine annealing mechanism to optimize the training process, thereby achieving superior performance on multiple public datasets.
Although mentioned SR methods have obtained excellent effects in SISR with certain scales, they may suffer from poor adaptability for real scenes. To address this problem, generative methods, i.e., flow-based models (Lugmayr et al., 2020b), VAEs (Heydari and Mehmood, 2020), diffusion model (Li et al., 2022a) are developed. Specifically, flow-based models directly calculate the mapping relationship between the normal distribution and the target distribution function to generate the target data (Jo et al., 2021). This processing brings both reversible data and huge computation cost. Variational Autoencoders (VAEs) extract hidden variable distributions through encoders, and then randomly sample them to generate new images (Chira et al., 2022). The training process of VAEs is stable, but the quality of generated images limits its application. Diffusion models generates images with outstanding diversity via iterating denoising (Gao et al., 2023b). However, it causes that the inference processing takes more time.
With the development of hardware devices, varieties of captured images are more(Li et al., 2025). However, scenes in the real world are varying, the number of obtained images are insufficient for real applications. To address this problem, generative methods, i.e., flow-based models, VAEs, diffusion model are developed. Specifically, flow-based models directly calculate the mapping relationship between the normal distribution and the target distribution function to better generate similar samples as given samples (Jo et al., 2021). Although it has better generative effect, it suffers from a challenge of huge computational cost. To solve unstable training of GANs(Gan et al., 2025), variational autoencoders (VAEs) used encoders to extract hidden variable distributions to overcome wave of training process (Chira et al., 2022). Although VAEs can be useful to accelerate training, they are limited by blurry image super-resolution. To enhance adaptive ability of obtained super-resolution models, diffusion models can use game strategies between adding and reducing noise to learn an unsupervised super-resolution model. Although a diffusion model can enhance robustness of obtained image super-resolution model, it may cause bigger computational costs and consume more resources. Taking into mentioned analysis account, GANs based game strategies are good tools of generative methods for image super-resolution.
@@
To address the problem of small samples, generative adversarial nets (GANs) used generator and discriminator in a game-like manner to obtain good performance on image applications (Goodfellow et al., 2014; Wang et al., 2021c). Specifically, the generator can generate new samples, according existing samples (Gui et al., 2021). The discriminator is used to distinguish the samples from generator (Gui et al., 2021). Due to their strong learning abilities, GANs become popular image super-resolution methods (Bell-Kligler et al., 2019). For instance, Park et al. (Park et al., 2023a) combined kernel idea and GAN to extract more structural information to enhance image sharpness and edge thickness for image super-resolution. However, there are few studies summarizing these GANs for SISR. Also, differing from previous work based deep learning techniques for image super-resolution, i.e., Refs.(Jiang et al., 2021; Anwar et al., 2020), we can not only refer to importance of GANs for low- and high-levels in terms of big and small samples, but also first deeply product a summary of GANs in image super-resolution, according to combination of different training ways (i.e., supervised, semi-supervised and unsupervised manners), network architectures, prior knowledge, loss functions and multiple tasks, which can makes readers easier know principle, improvements, superiority and inferiority of different GANs for image super-resolution. That is, in this paper, we conduct a comprehensive overview of 209 papers to show their performance, pros, cons, complexity, challenges and potential research points, etc. First, we show the effects of GANs for image applications. Second, we present popular architectures for GANs in big and small samples for image applications. Third, we analyze motivations, implementations and differences of GANs based optimization methods and discriminative learning for image super-resolution in terms of supervised, semi-supervised and unsupervised manners, where these GANs are worked by combining different network architectures, prior knowledge, loss functions and multiple tasks for image super-resolution. Fifth, we compare these GANs using experimental setting, quantitative analysis (i.e., PSNR, SSIM, complexity and running time) and qualitative analysis. Finally, we report on potential research points and existing challenges of GANs for image super-resolution. The overall architecture of this paper is shown in Fig 1.
The remainder of this survey is organized as follows: Section 2 reviews the developments of GANs; Section 3 surveys popular GANs for image applications; Section 4 focuses on introduction of existing GANs via three ways on SISR; Section 5 compares performance of mentioned GANs from Section 2 for SISR; Section 6 offers potential directions and challenges of GANs in image super-resolution; and Section 7 concludes the overview.
@@ 2. Developments of GANs
Traditional machine learning methods prefer to use prior knowledge to improve performance of image processing applications (Sun et al., 2010). For instance, Sun et al. (Sun et al., 2010) proposed a gradient profile to restore more detailed information for improving performance of image super-resolution. Although machine learning methods based prior knowledge has fast execution speed, they have some drawbacks. First, they required manual setting parameters to achieve better performance on image tasks. Second, they required complex optimization methods to find optimized parameters. According to mentioned challenges, deep learning methods are developed (Bird et al., 2022). Deep learning methods used deep networks, i.e., CNNs to automatically learn features rather than manual setting parameters to obtain effective effects in image processing tasks, i.e., image classification (Bird et al., 2022), image inpainting (Liu et al., 2021) and image super-resolution (Wang et al., 2020b). Although these methods are effective big samples, they are limited for image tasks with small samples (Goodfellow et al., 2014).
To address problems above, GANs are presented in image processing (Goodfellow et al., 2014). GANs consist of generator network and discriminator network. The generator network is used to generate new samples, according to given samples. The discriminator network is used to determine truth of obtained new samples. When generator and discriminator is balance, a GAN model is finished. The work process of GAN can be shown in Fig 2, where G and D denote a generator network and discriminator network. To better understand GANs, we introduce several basic GANs as follows.
@@
To obtain more realistic effects, conditional information is fused into a GAN (CGAN) to randomly generate images, which are closer to real images (Mirza and Osindero, 2014). CGAN improves GAN to obtain more robust data, which has an important reference value to GANs for computer vision applications. Subsequently, increasing the depth of GAN instead of the original multilayer perceptron in a CNN to improve expressive ability of GAN is developed for complex vision tasks (Radford et al., 2015). To mine more useful information, the bidirectional generative adversarial network (BiGAN) used dual encoders to collaborate a generator and discriminator to obtain richer information for improving performance in anomaly detection, which is shown in Fig 3(Donahue et al., 2016). In Fig 3, x denotes a feature vector, E is an encoder and y expresses an image from discriminator.
It is known that pretrained operations can be used to accelerate the training speed of CNNs for image recognition (Hinterstoisser et al., 2018). This idea can be treated as an energy drive. Inspired by that, Zhao et al. proposed an energy-based generative adversarial network (EBGAN) by using a pretraining operation into a discriminator to improve the performance in image recognition (Zhao et al., 2016). To keep consistency of obtained features with original images, cycle-consistent adversarial network (CycleGAN) relies on a cyclic architecture to achieve an excellent style transfer effect (Zhu et al., 2017) as illustrated in Fig 4.
@@
Although pretrained operations are useful for training efficiency of network models, they may suffer from mode collapse. To address this problem, Wasserstein GAN (WGAN) used weight clipping to enhance importance of Lipschitz constraint to improve the stability of training a GAN (Arjovsky et al., 2017). WGAN used weight clipping to perform well. However, it is easier to cause gradient vanishing or gradient exploding (Gulrajani et al., 2017). To resolve this issue, WGAN used a gradient penalty (treated as WGAN-GP) to break the limitation of Lipschitz for pursuing good performance in computer vision applications (Brock et al., 2018). To further improve results of image generation, GAN enlarged batch size and used truncation trick as well as BIGGAN can make a tradeoff between variety and fidelity (Brock et al., 2018). To better obtained features of different parts of an image (i.e., freckles and hair), style-based GAN (StyleGAN) uses feature decoupling to control different features and finish style transfer for image generation (Karras et al., 2019). The architecture of StyleGAN and its generator are shown in Fig 5 and Fig 6.
@@ 3. Popular GANs for image applications
According to mentioned illustrations, it is known that variants of GANs based on properties of vision tasks are developed in Section II. To further know GANs, we show different GANs on training data, i.e., big samples and small samples for different high- and low-level computer vision tasks as shown in Fig 7.
@@ 3.1.1. GANs on big samples for image generation
Good performance of image generation depends on rich samples. Inspired by that, GANs are improved for image generation (Gui et al., 2021). That is, GANs use generator to produce more samples from high-dimensional data to cooperate discriminator for promoting results of image generation. For instance, boundary equilibrium generative adversarial networks (BEGAN) used obtained loss from Wasserstein to match loss of auto-encoder in the discriminator and achieve a balance between a generator and discriminator, which can obtain more texture information than that of common GANs in image generation (Berthelot et al., 2017). To control different parts of a face, StyleGAN decoupled different features to form a feature space for finishing transfer of texture information (Karras et al., 2019). Enhanced GAN for image generation (EIGGAN) (Bergmann et al., 2017) improves the performance of the generator by incorporating a spatial attention mechanism and parallel residual operations, thereby achieving higher quality and more realistic image generation effects in large-scale image generation tasks. Besides, texture synthesis is another important application of image generation (Li and Wand, 2016). For instance, Markovian GANs (MGAN) can quickly capture texture date of Markovian patches to achieve function of real-time texture synthesis (Li and Wand, 2016), where Markovian patches can be obtained Ref. (Gui et al., 2021). Periodic spatial GAN (PSGAN) (Bergmann et al., 2017) is a variant of spatial GAN (SGAN) (Jetchev et al., 2016), which can learn periodic textures of big datasets and a single image. These methods can be summarized in Table 2.
Figure 7. Frame of popular GANs for image applications.
@@
EIGGAN (Tian et al., 2024a) GAN GAN with spatial attention for image generation